Title Short form of title for page headings Author’s name (or authors’ names) Short form of names First address line

Second address line Third address line ... a.b@x.y.z

Abstract

Put the abstract here.

1 Section heading

1.1 Subsection heading

1.1.1 Subsubsection heading

Here is an example.

Here is a definition. Here is a lemma. Here is a theorem. Proof using the proof environment. Here is a conjecture. Here is a remark. The full list of environments defined by the style file is: blank - a new

numbered paragraph, claim - Claim, cond - Conditions, conj - Conjecture cor - Corollary, Def - Definition, diag - Diagram, lemma - Lemma prop - Proposition, qn - Question, result - Result, and theorem - Theorem.

Also there are unnumbered environments: aside - Comment, notn - Notation, proof - Proof and rem - Remark.

Some maths commands are already defined, including , and for the reals, integers and natural numbers.

Please use italics when defining a word for the first time. Put any acknowledgements here.

References

[1] names of author(s) paper title journal title volume year 2000–3000

[2] name of author(s) (should be A. N. Onymous format) article title journal title book title proceedings title editor name, single editor series name for book series publisherlocation of publisher various extra details volume number for journals year of journal for year of book for year of proceedings page numbers page numbers in a proceedings